OpenStudy (anonymous):

Solve: e^2sinx=1, for 0≤x≤4

OpenStudy (freckles):

recall \[e^0=1 \\ \text{ so solve } 2\sin(x)=0 \text{ on } 0 \le x \le 4\]

OpenStudy (freckles):

also I assumed the problem was \[e^{2 \sin(x)}=1\]

OpenStudy (anonymous):

Yes :3

OpenStudy (freckles):

let me know if you need any further help

OpenStudy (anonymous):

I looked online, and it said the answer was x=0±2πn,π±2πn But i don't understand where the 2πn came from.

OpenStudy (freckles):

sin(x) has period 2pi

OpenStudy (anonymous):

Oh yea! Haha. sorry. It's been a while...

OpenStudy (freckles):

so sin(x)=sin(x+2pi)=sin(x+4pi)=sin(x+6pi)=... or sin(x)=sin(x+2*npi) where n is an integer

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

Since \(0 \le x \le 4\), doesn't that restrict the solutions to only a few values of x, and not x=0±2πn,π±2πn in general for any integer n. The solutions are only x = 0, π

OpenStudy (anonymous):

Oh, i get it better now! Thanks!

OpenStudy (mathstudent55):

You're welcome.